Circular No. 6670

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 G. Pooley and D. Green, Mullard Radio Astronomy Observatory, Uni-

versity of Cambridge, report: “We obtained observations of the radio source (IAUC 6662) associated with GRB 970508 with the Ryle Telescope at 15 GHz on seven successive days. During May 16.45–16.58 UT the mean flux density was 1.57 ± 0.25 mJy (1σ), a result supporting the inverted radio spectrum reported by Frail et al. (IAUC 6662). On none of the other six days was an individually significant detection made, but by averaging the data obtained during May 17–22 we obtain a measurement significant at the 3σ level, giving a flux density of 0.66 ± 0.11 mJy (1σ).”

G. B. Taylor, A. J. Beasley and D. A. Frail, National Radio Astronomy Observatory; and S. R. Kulkarni, California Institute of Technology, report: “Observations with the Very Long Baseline Array at 8.4 GHz have detected the radio source (IAUC 6662) identified with Bond’s optical variable (IAUC 6654, 6655) within the error circle of GRB 970508. During May 17.17–17.40 UT an unresolved source (size < 0′′.0003) was found at α = 6^{h}53^{m}49^{s}.4513, δ = +79^{o}16′19′′.513 (equinox J2000.0). This position is consistent with the earlier VLA position (IAUC 6662, with errors of ±0′′.1), but the current uncertainty in the position derived from the VLBA observations is 0′′.001 in both coordinates. The flux density of the source was 0.48 ± 0.06 mJy. Phase referencing was performed by cycling every 3 min between the target source and the strong (0.75 Jy) compact source B0718+793 located 1^{o}.5 away. A preliminary reduction of observations during May 16.48–16.92 using the VLBA, the 100-m radio telescope at Effelsberg and the phased VLA confirms this detection and should yield an image with much lower thermal noise. We are continuing to observe this source every few days with the VLBA.”

COMET P/1997 G1 (MONTANI) As was suggested on IAUC 6624, further observations through May 19

have shown this to be a short-period comet. The following orbital elements are taken from MPC 29880:

T = 1997 Apr. 26.757 TT  ω = 216.^{o}527
e = 0.45668  Ω = 267.471 }2000.0
q = 4.24044 AU  i = 3.959
a = 7.80461 AU n^{o} = 0.045204 P = 21.80 years
1997 May 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden